# Generating Temporal Networks with the Ascona Model

### v1

@@ 4 Conclusions
In Section˜3.3, we introduced smooth SBM, a specific variation of the Ascona model that generates link streams with community structures, and compared it with the dynamic SBM model. Similarly, one could design alternative (conditional) probability connectivity matrices to generate structures with different properties, e.g., nestedness, disassortativity, and structurelessness. In Section˜3.4, we have seen how to generate archetypical temporal structures with the smooth SBMs. As discussed in [failla_describing_2024], hybridization of these simple archetypes leads to more complex structures. It is realistic to imagine generating hybrids by combining several queue blocks, and this is a possible future direction of this work, raising the question of how expressive the Ascona model really is.
Aside from these positive aspects of expressivity and flexibility, a big advantage of this article’s approach is that it translates temporal graphs’ problems to problems of time series and stochastic processes that are well studied and understood, with a rich body of theoretical and practical results ready to be exploited. For instance, [walrand_introduction_1988] discusses various aspects of statistical analysis on queues. $M/M/\infty$ models, the foundation of the EDEDE structures, are particularly nicely behaved from a statistical perspective: for instance, the MLE estimation is equivalent to the method of moments when fitting the exponential distributions of the model. Other queue models would require the use of simulations, but some are remarkably amenable to analyze (e.g., see [eick_physics_1993]).
Future work should address aspects of inferential statistics: given a reference link stream, find the Ascona model parametrization that fits the data in the best way possible. Recent works on maximum entropy approaches to temporal networks [barucca_maximum_2025, clemente_temporal_2024] could pave the way in this direction and go beyond the purely Markovian framework.
@@
A GitHub project containing the code used for the generation of all plots is available [^2] https://github.com/samuelkoovely/EDEDE_project.
@@ Appendix A Head of the Queue: Computations
$\rho\coloneqq\frac{\lambda}{\mu},\qquad x\coloneqq e^{-\mu t^{\star}},\quad\text{so that}\quad t^{\star}=-\frac{\ln x}{\mu}.$ (A.2)
@@
Because $x=e^{-t^{\star}/\mu}$ must be positive, we take the ‘++’ root:
@@ Acknowledgments
Y. Asgari and G. Vaccario suggested articles relevant to the subject, while A. Bovet and Z. Fang’s comments helped me to improve this manuscript. C. Casati helped in finding the catchy palindromic acronym EDEDE. My gratitude goes to all of them.

### v2

@@ 4 Conclusions
In Section˜3.2.2, we introduced smooth SBM, a specific variation of the Ascona model that generates link streams with community structures, and compared it with the dynamic SBM model. Similarly, one could design alternative (conditional) probability connectivity matrices to generate structures with different properties, e.g., nestedness, disassortativity, and structurelessness. In Section˜3.2.3, we have seen how to generate archetypical temporal structures with the smooth SBMs. As discussed in [14], hybridization of these simple archetypes leads to more complex structures. It is realistic to imagine generating hybrids by combining several queue blocks, and this is a possible future direction of this work, raising the question of how expressive the Ascona model really is.
Aside from these positive aspects of expressivity and flexibility, a big advantage of this article’s approach is that it translates temporal graphs’ problems to problems of time series and stochastic processes that are well studied and understood, with a rich body of theoretical and practical results ready to be exploited. For instance, [34] discusses various aspects of statistical analysis on queues. $M/M/\infty$ models, the foundation of the EDLDE structures, are particularly nicely behaved from a statistical perspective: for instance, the MLE estimation is equivalent to the method of moments when fitting the exponential distributions of the model. Other queue models would require the use of simulations, but some are remarkably amenable to analyze (e.g., see [13]).
Future work should address aspects of inferential statistics: given a reference link stream, find the Ascona model parametrization that fits the data in the best way possible. Recent works on maximum entropy approaches to temporal networks [4, 11] could pave the way in this direction and go beyond the purely Markovian framework.
@@
A GitHub project containing the code used for the generation of all plots is available [^2] https://github.com/samuelkoovely/EDLDE_project.
@@ Appendix A Head of the Queue: Computations
$\rho\coloneqq\frac{\lambda}{\mu},\qquad x\coloneqq e^{-\mu t^{\star}},\quad\text{so that}\quad t^{\star}=-\frac{\ln(x)}{\mu}.$ (A.2)
@@
Because $x=e^{-\mu t^{\star}}$ must be positive, we take the ‘++’ root:
@@ Acknowledgments
Y. Asgari and G. Vaccario suggested articles relevant to the subject, while A. Bovet and Z. Fang’s comments helped me to improve this manuscript. C. Casati helped in finding the catchy palindromic acronym EDLDE. My gratitude goes to all of them.